What is the pH of a 0.250 M aqueous solution of ammonium chloride (NH4+Cl-)?                              &nb

  1. What is the pH of a 0.250 M aqueous solution of ammonium chloride (NH4+Cl-)?

                                                                                                            Kb (NH3) = 1.2 × 10-5

    A.

    5.04

    B.

    4.95

    C.

    4.77

    D.

    4.84

    E.

    6.26
Answer #1

NH4Cl is weak acid because it gets dissociated in aqueous medium in following ways,

NH4Cl(s) + NH(aq) + C1-(aq)

NH(aq) + H2O(1) = H2O+ (aq) + NH4OH(aq)

So, at equilibrium, in solution of weak acid, as we know,

10-14 [H30+] = VKq xC= v kaxo

= [H3O+] = V 10-14 V 1.2 x 10-5 X 0.25010-14 = [H3O+]=V 1.2 x 10-5 X 0.250 = 1.44337 x 10-5

Thus, pH = -log[H3O+] = -log(1.44337 x 10-5) = 4.84

So, correct answer is D.